Find the Tangent and Normal line to the given curve. y=√(16+x^2) at the origin.



1
Expert's answer
2021-01-11T14:33:35-0500

Solution


  • At x=0 y=4 means that the curve does not pass through the origin.
  • When the derivative is considered to estimate the turning points,

"\frac{dy}{dx}=\frac{x}{\sqrt{16+x^2}}=0\\
x=0"

  • This means that the curve has a turning point at x=0 simultaneously y = 4 at that point.
  • As x reaches infinity y reaches infinity concluding that this curve is has a minimum at (0,4) expands towards y's positive direction.
  • Therefore, a tangent or a normal could be drawn at (0,4) not at the origin.
  • And as it is easy to figure out, the tangent drawn at (0,4) has a gradient of 0 (according to the derivative at x=0) hence the tangent is parallel to x axis. Since y=4 at x=0, line parallel to x axis at this point is y=4
  • Again any normal to the x axis is parallel to the y axis hence at that point, the normal is x=0 turning out to be again the y axis.


  • Tangent: y=4
  • Normal: x=0/y axis